OpenStudy (anonymous):

Francisco is playing a game with 3 green, 2 yellow, 4 red, and 3 black marbles in a bag. He has calculated the probability of drawing a yellow marble, not replacing it, and then drawing a red marble. Explain the error in his solution. (2/12)(4/12)=8/144

OpenStudy (anonymous):

Can anyone help..... anyone?

OpenStudy (anonymous):

not replacing it?

OpenStudy (anonymous):

not replacing it.

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

As far as I can understand the language, the yellow marble gets drawn when he first seects it so the number of marbles now become one less, i.e 12 - 1 = 11 correct answer should be 2/12 (4/11) = 8/132

OpenStudy (anonymous):

Thanks you are correct once more!